a proton has total energy of 100 Gev what is its approximate velocity?
Answer #1


0.5*m*v^2 = 100GeV.....

speed v = sqrt((100*10^9*1.6*10^-19)/(0.5*1.67*10^-27)) = 4.377*10^9 m/sec